can u help me solvethese 3 problems?
1.x times1 over y?
2.what is a 16oz of bottle of water costs $1.44.what is the cost per ounce?
3.Evaluate 3(x^3-5x)+6for x=3?

Answer by AnlytcPhil(1807)

can u help me solvethese 3 problems?
1.  x times 1 over y?

     1     x     1     1·x      x
x · ——— = ——— · ——— = —————— = ———
     y     1     y     y·1      y

2.what is a 16oz of bottle of water costs $1.44.what is the cost per ounce? 

1.44 dollars    144 cents     144   cents      cents
———————————— = ———————————  = ——— · —————— = 9 ————— = 9¢/oz
  16 ounces      16 ounces     16   ounce      ounce


3.Evaluate 3(x3-5x)+6for x=3?

     3(x3 - 5x) + 6

Change the parentheses "{ }" to brackets "[ ]"

     3[x3 - 5x] + 6

In place of each "x" write "(3)"

     3[(3)3 - 5(3)] + 6

(3)3 = 3·3·3 = 27 so replace the (3)3 by 27

     3[27 - 5(3)] + 6

5(3) = 15, so replace the 5(3) by 15

     3[27 - 15] + 6

27 - 15 = 12 so replace the 27 - 15 by 12

     3[12] + 6

3[12] = 36 so replace the 3[12] by 36

      36 + 6

36 + 6 = 42, so replace the 36 + 6 by 42

      42 

Edwin
AnlytcPhil@aol.com